I need your help thank you!
Modern Algebra Problems:
The Matriarch.
"You ash how old I am, doctor? Well, I'll tell you! I was a girl for the first sixth of my life; then I studied at the University for the next twelfth. Upon graduation, I met the man who married me three years later; and we devoted the next third of my life to raising our children. How quickly they grew, and how soon they left! After one-fourth more of my life had passed me by, our first great-grandchild was born. I doted on my descendants for the next eighth of my life. Last year my beloved husband died; soon I will join him. Looking back, I regret nothing.

Answer by mananth(16946)

Well, I'll tell you!
let my age be x years

I was a girl for the first sixth of my life; x/6...................16
then I studied at the University for the next twelfth. x/12...............8
Upon graduation, I met the man who married me three years later;3 years...3
and we devoted the next third of my life to raising our children. x/3.....32
After one-fourth more of my life had passed me by, x/4................24
I doted on my descendants for the next eighth of my life. x/8...........12

Last year my beloved husband died; +1...................................1
x/6 + x/12 + x/3 + x/4 +x/8 + 4 =x
(8x+4x+16x+12x+6x)/48 +4=x
x-46/48 x =4
2x/48 =4
2x=192
x=96